OpenStudy (skittles_for_life6422):

!!!!!!!!!!!!!!!!!!!!!!!!!WILL GIVE MEDAL AND FAN!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!HELP!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!

OpenStudy (skittles_for_life6422):

\[-11\sqrt{112}\]

OpenStudy (jdoe0001):

is that a statement or a question?

OpenStudy (skittles_for_life6422):

ANSWER CHOICES \[-44\sqrt{7}, -176\sqrt{7}, -27\sqrt{7}, 4\sqrt{7}\]

OpenStudy (skittles_for_life6422):

QUESTION @jdoe0001

OpenStudy (skittles_for_life6422):

sorry that was in caps

OpenStudy (jdoe0001):

question? I don't see it there it's just a value with a radical and a radicand

OpenStudy (jdoe0001):

not more of a question than say "2" or 53

OpenStudy (skittles_for_life6422):

oh whoops you have to simplify

OpenStudy (solomonzelman):

\(\color{black}{ \displaystyle -11\sqrt{112} }\) how I visualized it: \(\color{black}{ \displaystyle 112=4\times (25+3) =4\times 28=4\times 4 \times 7 }\) your hint from there: \(\color{black}{ \displaystyle 112=4^2\times 7 }\)

OpenStudy (skittles_for_life6422):

?

OpenStudy (jdoe0001):

\(\large \bf { -11\sqrt{112} \\ \quad \\ {\color{brown}{ 112\to 2\cdot 2\cdot 2\cdot 2\cdot 7\to 2^2\cdot 2^2\cdot 7\to 4^2\cdot 7}} \\ \quad \\ -11\sqrt[2]{{\color{brown}{ 4^2\cdot 7 }}} }\)

OpenStudy (skittles_for_life6422):

? @SolomonZelman

OpenStudy (solomonzelman):

yes, jdoee0001 posted it:)

OpenStudy (skittles_for_life6422):

thats is much clearer

OpenStudy (solomonzelman):

yes, much more clear indeed..... so what do you get ?

OpenStudy (skittles_for_life6422):

\[-44\sqrt{7}\]

OpenStudy (skittles_for_life6422):

@SolomonZelman

OpenStudy (skittles_for_life6422):

@jdoe0001

OpenStudy (jdoe0001):

yeap

OpenStudy (solomonzelman):

yes, correct !

OpenStudy (skittles_for_life6422):

Thanks guys u rock @SolomonZelman @jdoe0001

OpenStudy (jdoe0001):

yw

OpenStudy (skittles_for_life6422):

I forgot can i ask my last question @SolomonZelman @jdoe0001

OpenStudy (skittles_for_life6422):

i will open a new question though